A 25- Liter solution of alcohol in water is 72% alcohol. How many liters of an 80% solution must be added if the final solution should be 75% alcohol?.

Answer by mananth(16946)

-------------- percent ---------------- Amount
Alcohol I 72 ---------------- 25 liters
Alcohol II 80 ---------------- x liters
Mixture 75 ---------------- 25 + x liters

72*25+80x=75(25 +x)
1800+80x=1875+75x
80x-75x =-1800+1875
5 x = 75
/ 5
x = 15 liters Alcohol II

m.ananth@hotmail.ca